Homework (Ch 08) Attempts: Average: /5 4. The Laffer curve Governments often place so-called sin taxes on goods or services such as cigarettes, alcohol, and pornography. These kinds of taxes are popular with politicians because they are usually more palatable to voters than income taxes. To understand the effect of such a tax, consider the monthly market for adult DVDs, which is shown on the following graph Use the graph input tool to help you answer the following questions. You will not be graded on any changes you make to this graph Note: Once you enter a value in a white field, the graph and any corresponding amounts in each grey field will change accordingly. Graph Input Tool Market for Adult DVDs 40 36 32 28 24 Quantity DVDs) 12 ermianeriv36.00a PecvD) 4.00 (Dollars per DVD) Tax Wedge (Dollars per DVD) 32.00 e 16 12 0 12 24 3 48 60 72 84 96 108 120 QUANTITY (DVDs) Suppose the government imposes an $8-per-DVD tax on suppliers. At this tax amount, the equilibrium quantity of adult DVDs is 48 DVDs, and the government collects $384 in tax revenue Now calculate the governments tax revenue if it sets a tax of $0, $8, $16, $20, $24, $32, or $40 per DVD. (Hint: To find the equilibrium quantity after the tax, adjust the Quantity field until the Tax Wedge equals the value of the per-unit tax.) Using the data you generate, plot a Laffer curve by using the green points (triangle symbol) to plot total tax revenue at each of those tax levels.Homework (Ch 08) using the green points (triangle symbol) to plot total tax revenue at each of those tax levels Note: Plot your points in the order in which you would like them connected. Line segments will connect the points automatically 960 884 Laffer Curve 788 672 578 480 384 288 96 04 18 20 24 28 32 3 40 TAX (Dollars per DVD) Suppose the government is currently imposing a $24-per-DVD tax on adult DVDs True or False: The government can raise its tax revenue by decreasing the per-unit tax on adult DVDs. True O False Consider the deadweight loss generated in each of the following cases: no tax, a tax of $16 per DVD, and a tax of $32 per DVD. On the following graph, use the black curve (plus symbols) to illustrate the deadweight loss in these cases. (Hint: Remember that the area ofa triangle is equal to 흘 x Basex eight. In the case of a deadweight loss triangle found on the graph input too, the base is the amount of the tax and the height is the reduction in quantity caused by the tax.)Consider the deadweight loss generated in each of the following cases: no tax, a tax of $16 per DVD, and a tax of $32 per DVD On the following graph, use the black curve (plus symbols) to illustrate the deadweight loss in these cases. (Hint: Remember that the area of a triangle is equal to 흘 × Base × Height. In the case of a deadweight loss triangle found on the graph input tool, the base is the amount of the tax and the height is the reduction in quantity caused by the tax.) 864 Deadweight Loss 788 672 578 480 384 288 192 86 4 8 12 20 24 28 32 38 40 TAX (Dollars per DVD) As the tax per DVD increases, deadweight loss increases by a greater and greater amount Grade It Now Save & Continue Continue without saving
Answer #1

The governemnt is imposing a tax of $8 on DVDs which means the wedge on the graph would be when the difference between the prices at which the CDs are demanded and supplied is $8 .

140 36 32 28 IG -T-t-t-一ㄒㄧㄒㄧㄒㄧㄒㄧ t 12-24 36 48 6072 84 %108120 o - (DVDS

At this the government collects a tax of $8 and the equilibrium level of quantity is 48 and the revenue collected by government is $48*8= $384

The table for laffer curve is as given

Tax DVDs (quantity ) Revenue
0 60 0
8 48 384
16 36 576
20 30 600
24 24 576
32 12 384
40 0 0

(14,576) (20,600 04,576) 420 心 384 (32, 384 2.8 8 192 9 6 8 12 20 24 28 32 36 10 Tat

These points are marked on laffer curve

When the tax is $24 , then the government can increase the tax revenue by decreasing the tax rate.

The statement is true because as in the table the at the tax rate $24 , the revenue is $576 and when the tax rate is decreased ie at 20 , the tax revenue is $600.

Deadweight loss at no tax is none.

Deadweight loss at $16 tax is given by the area of triangle between new and old equilibrium

Area of deadweight tax = 1/2*b*h = 1/2 *(tax )*( old equil- new equil)

= 1/2 *16*(60-36)

= 192 sq units

Area of deadweight tax at $32 = 1/2*b*h = 1/2 *(tax )*( old equil- new equil)

= 1/2*32*(60-12)

= 768

768 622 F 521 (32768) d576 9 80 384 288 192 46 0 4 2 l6 20 24 28 32 36 4 0

As the tax per DVD increases,   deadweight losses increases by greater and greater amount.

(You can comment for doubts )